A 10-kg box is sitting initially on a rough flat surface. The static and kinetic coefficient of friction of the surface and the box is 0.32 and 0.2, respectively. What is the magnitude of the friction force acting on the box while the force of 25 N is applied on it?
1
Expert's answer
2018-11-13T12:59:11-0500

mass:


m=10kg

static coefficient of friction :

µ_s=0.32

dinamic coefficient of friction :

µ_d=0.2

Force:

F=25N

Solution:

static force of friction :

F_s=N*µ_s=mgµ_s=31.4N


dinamic force of friction :

F_d=N*µ_d=mgµ_d=19.6N


Answer:

static force of friction :

31.4N


dinamic force of friction :

19.6N